OpenStudy (anonymous):

Two investments totaling $64,000 produce an annual income of $3410. One investment yields 2% a year while the other yields 8% per year. How much is invested at each rate?

OpenStudy (amistre64):

3410/64000 = some convoluted percentage .... a + b = t ; a = t-b a(x) + b(y) = i (t-b)(x) + b(y) = i tx + b(y-x) = i b at y% = (i-tx)/(y-x)

OpenStudy (amistre64):

b at 8% = (3410-64000(.02))/(.08-.02) .. let me dbl chk that tho

OpenStudy (amistre64):

35500(.08) + 28500(.02) = 3410

OpenStudy (amistre64):

it worked :)

OpenStudy (anonymous):

so 3410 is invested at what percent?

OpenStudy (amistre64):

3410 is how much money was made off of the investments in total; so: 35500 at 8% = 3410 28500 at 2%